# Characteristic Learning for Provable One Step Generation

### v3

@@ Appendix H Approximation by Deep Neural Networks with Lipschitz Constraint
$A_{k}(m)=\prod_{\ell\neq k}\psi\Big{(}3N_{\ell}\Big{(}x_{\ell}-\frac{m_{\ell}}{N_{\ell}}\Big{)}\Big{)}\mathbbm{1}\Big{\{}m:\Big{|}x_{\ell}-\frac{m_{\ell}}{N_{\ell}}\Big{|}\leq\frac{2}{3N_{\ell}},\leavevmode\nobreak\ \ell\neq k\Big{\}}.$
@@
We next estimate (H.7) in the following cases.
@@
We have investigated the approximation error of a target function on the hypercube $[0,1]^{p}$ in Lemma H.5. In the following corollary, we extend our analysis to target functions on general bounded domain $[0,T]\times[-R,R]^{d}$.
Corollary H.6.
@@
Set the deep neural network class $N(L,S)$ as $L=\lceil\log_{2}p\rceil+3$ and $S=(22p+6)\prod_{k=1}^{p}(N_{k}+1)$. According to Lemma H.5, there exists a neural network $u\in N(L,S)$ such that
@@
(H.8) $\leq 2^{p}\sum_{k=1}^{p}\frac{b_{k}-a_{k}}{N_{k}}\mathop{\operatorname{ess}\,\sup}_{x\in{\mathcal{X}}}\Big{|}\frac{\partial u^{*}}{\partial x_{k}}(x)\Big{|},$
@@
(H.9) $\mathop{\operatorname{ess}\,\sup}_{x^{\prime}\in[0,1]^{p}}\Big{|}\frac{\partial u}{\partial x_{k}^{\prime}}(x^{\prime})\Big{|}\leq 3\mathop{\operatorname{ess}\,\sup}_{x^{\prime}\in[0,1]^{p}}\Big{|}\frac{\partial(\phi\circ u^{*})}{\partial x_{k}^{\prime}}(x^{\prime})\Big{|}=3(b_{k}-a_{k})\mathop{\operatorname{ess}\,\sup}_{x\in{\mathcal{X}}}\Big{|}\frac{\partial u^{*}}{\partial x_{k}}(x)\Big{|}.$
@@
where $(x^{\prime}-a)/(b-a)=((x_{k}^{\prime}-a_{k})/(b_{k}-a_{k}))_{k=1}^{p}\in{\mathbb{R}}^{p}$. It follows from (H.9) that
@@
for each $1\leq k\leq p$. Then composing $\psi$ on both sides of (H.8) yields the desired inequality. ∎
Appendix I Denoiser parameterization
@@
(I.1) $D_{\theta}(t,x)=c_{\text{skip}}(t)x+c_{\text{out}}(t)F_{\theta}\left(c_{\text{noise}}(t),c_{\text{in}}(t)x\right),$
@@
(I.2) ${\mathcal{L}}(F)=$ $\int_{0}^{1}{\mathbb{E}}_{X_{0}}{\mathbb{E}}_{X_{1}}\left[\omega(t)\|c_{\text{skip}}(t)X_{t}+c_{\text{out}}(t)F\left(c_{\text{noise}}(t),c_{\text{in}}(t)X_{t}\right)-x_{1}\|^{2}\right]\,\mathrm{d}t$
@@
(I.3) $\Leftrightarrow$ $c_{\text{out}}(t)=\sqrt{(1-\alpha_{t}c_{\text{skip}}(t))^{2}\sigma^{2}+c_{\text{skip}}(t)^{2}\beta_{t}^{2}}.$
@@
(I.4) $\Leftrightarrow$ $c_{\text{skip}}(t)=\frac{\alpha_{t}\sigma^{2}}{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}.$
We can check that (I.4) is indeed the minima of $c_{\text{out}}$. Meanwhile, (I.3) yields
@@
$c_{\text{noise}}(t)$ free choice
@@
Now (I.2) can be used as the working denoiser matching loss.
Appendix J Extra experiment details
@@
We use DDPM++ in as the denoiser network architecture, and embed another temporal input into it to construct the characteristic generator as Kim et al. [2024] did. We use Rectified Adam (RAdam) as the optimizer and set learning rate to 0.001 on MNIST and 0.0001 on CIFAR-10. In practice, we initialize $D_{{\mathcal{S}}}$ from the pre-trained $D_{{\mathcal{T}}}$ for fast convergences. Inspired by Esser et al. [2021], Kim et al. [2024], we adaptively balance the the characteristic matching loss and the denoiser matching loss. For characteristic matching on image data, we use the Learned Perceptual Image Patch Similarity (LPIPS) metric instead of $L^{2}$-norm to measure distance [Zhang et al., 2018]. We use exponential moving average (EMA) with rate 0.999 on MNIST and 0.9999 on CIFAR-10. For the choice of $u$ in (4.8), we apply a weighted random strategy which is more likely to choose a long range interval $[t,u]$ to ensure the precision of the teacher solver, as Kim et al. [2024] did, For each batch, we run the teacher solver with at most 19 NFE. On MNIST, we calculate FID in original pixel space. On CIFAR-10, we use Inceptionv3 to extract features (dimension=2048) of images first and then calculate FID in the feature space.

### v5

@@ Appendix I Generalization Error Analysis for Least-Squares Regression
$A_{k}(m)=\prod_{\ell\neq k}\psi\Big(3N_{\ell}\Big(x_{\ell}-\frac{m_{\ell}}{N_{\ell}}\Big)\Big)\mathbbm{1}\Big\{m:\Big|x_{\ell}-\frac{m_{\ell}}{N_{\ell}}\Big|\leq\frac{2}{3N_{\ell}},\penalty 10000\ \ell\neq k\Big\}.$
@@
We next estimate (J.7) in the following cases.
@@
We have investigated the approximation error of a target function on the hypercube $0,1]^{p}$ in Lemma J.5. In the following corollary, we extend our analysis to target functions on general bounded domain $0,T]\times[-R,R]^{d}$.
Corollary J.6.
@@
Set the deep neural network class $N(L,S)$ as $L=\lceil\log_{2}p\rceil+3$ and $S=(22p+6)\prod_{k=1}^{p}(N_{k}+1)$. According to Lemma J.5, there exists a neural network $u\in N(L,S)$ such that
@@
(J.8) $\leq 2^{p}\sum_{k=1}^{p}\frac{b_{k}-a_{k}}{N_{k}}\mathop{\operatorname{ess}\,\sup}_{x\in{\mathcal{X}}}\Big|\frac{\partial u^{*}}{\partial x_{k}}(x)\Big|,$
@@
(J.9) $\mathop{\operatorname{ess}\,\sup}_{x^{\prime}\in[0,1]^{p}}\Big|\frac{\partial u}{\partial x_{k}^{\prime}}(x^{\prime})\Big|\leq 3\mathop{\operatorname{ess}\,\sup}_{x^{\prime}\in[0,1]^{p}}\Big|\frac{\partial(\phi\circ u^{*})}{\partial x_{k}^{\prime}}(x^{\prime})\Big|=3(b_{k}-a_{k})\mathop{\operatorname{ess}\,\sup}_{x\in{\mathcal{X}}}\Big|\frac{\partial u^{*}}{\partial x_{k}}(x)\Big|.$
@@
where $(x^{\prime}-a)/(b-a)=((x_{k}^{\prime}-a_{k})/(b_{k}-a_{k}))_{k=1}^{p}\in{\mathbb{R}}^{p}$. It follows from (J.9) that
@@
for each $1\leq k\leq p$. Then composing $\psi$ on both sides of (J.8) yields the desired inequality. ∎
Appendix K Denoiser parameterization
@@
(K.1) $D_{\theta}(t,x)=c_{\text{skip}}(t)x+c_{\text{out}}(t)F_{\theta}\left(c_{\text{noise}}(t),c_{\text{in}}(t)x\right),$
@@
(K.2) ${\mathcal{L}}(F)=$ $\int_{0}^{1}{\mathbb{E}}_{X_{0}}{\mathbb{E}}_{X_{1}}\left[\omega(t)\|c_{\text{skip}}(t)X_{t}+c_{\text{out}}(t)F\left(c_{\text{noise}}(t),c_{\text{in}}(t)X_{t}\right)-x_{1}\|^{2}\right]\,\mathrm{d}t$
@@
(K.3) $\Leftrightarrow$ $c_{\text{out}}(t)=\sqrt{(1-\alpha_{t}c_{\text{skip}}(t))^{2}\sigma^{2}+c_{\text{skip}}(t)^{2}\beta_{t}^{2}}.$
@@
(K.4) $\Leftrightarrow$ $c_{\text{skip}}(t)=\frac{\alpha_{t}\sigma^{2}}{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}.$
We can check that (K.4) is indeed the minima of $c_{\text{out}}$. Meanwhile, (K.3) yields
@@
$c_{\text{noise}}(t)$ - free choice
@@
Now (K.2) can be used as the working denoiser matching loss.
Appendix L Extra experiment details
This section reports the detailed settings of the experiments conducted on image datasets. The details are shown in Table 9.
Table 9. Hyperparameters.
MNIST CIFAR-10 CelebAHQ-256 CelebAHQ-512
encoder/decoder
number of GPUs 4 $\times$ A800 4 $\times$ A800 12 $\times$ A100 8 $\times$ A100
architecture DDPM++ DDPM++ DiT-B/2 DiT-XL/2
channel multiplier 32 128 - -
channels per resolution 1, 2, 2 1, 2, 2, 2 - -
model size (GB) 0.02 0.5 0.25 1.3
optimizer RAdam RAdam AdamW AdamW
learning rate 1e-3 1e-4 1e-4 1e-4
LR ramp-up (Mimg) 0.5 0.5 10 10
EMA rate 0.999 0.9999 - -
EMA half-life (Mimg) - - 0.5 0.5
teacher duration (Mimg) 10 100 - -
student duration (Mimg) 2 2 100 15
batch size 1024 1024 192 128
stochastic interploants Föllmer Föllmer Linear Linear
global loss metric LPIPS  [^1] Zhang et al. [2018] LPIPS $L^{2}$ $L^{2}$
adaptive weightning  [^2] Esser et al. [2021], Kim et al. [2024]
NFE of teacher (max) 19 19 2 2
dropout 0.1 0.1 0.13 0.13
@@
In practice, when there is an available teacher network $D_{{\mathcal{T}}}$, we initialize $D_{{\mathcal{S}}}$ from the pre-trained $D_{{\mathcal{T}}}$, and then embed another temporal input into it to construct the characteristic generator as Kim et al. [2024] did. For the choice of $u$ in (4.8), we apply a weighted random strategy which is more likely to choose a long range interval $[t,u]$ to ensure the precision of the teacher solver, as Kim et al. [2024] did. Inspired by Esser et al. [2021], Kim et al. [2024], we adaptively balance the global characteristic matching loss and the local denoiser matching loss on MNIST and CIFAR-10. We calculate FID in original pixel space for MNIST, and in the feature space (dimension=2048) extracted by a pre-trained Inceptionv3 network for other image datasets.
Appendix M Remarks on Time Consumption
For numerical experiments on the MNIST and CIFAR-10 dataset, a pre-trained denoiser network (teacher model) is explicitly involved, to generate the reference trajectories, from which the characteristic generator is distilled.
Take the CIFAR-10 dataset for example. The total time consumption is summarized as follows: we first spend around 40 NVIDIA A800 hours during the pre-train stage of the denoiser network. During the training of the characteristic generator, the maximum number of steps to generate each reference trajectory is set to 19. On average, the reference trajectory is generated on-the-fly with speed 0.007 sec per image. In total, the training time of the characteristic generator is around 3 NVIDIA A800 hours, 38% of which is spent on generating the reference trajectory from the teacher model.
For the CelebA 256 $\times$ 256 dataset, we first spend around 78 NVIDIA A800 hours during the pre-train stage of the denoiser network. Then we use the pre-trained result from the first stage as the source of parameter initialization and teacher model. The training time of characteristic generator is around 480 NVIDIA A800 hours, around 45% of which are spent on generating a reference trajectory, by evaluating the characteristic generator itself twice.
For the CelebA 512 $\times$ 512 dataset, there is no pre-training stage. The training time of characteristic generator is around 480 NVIDIA A800 hours.
It is important to note that these training times are one-time offline costs. Once trained, the characteristic generator can significantly reduce the number of function evaluations (NFE) during inference. As demonstrated in Table 5 of Section 4.2, traditional sampling methods for diffusion models require hundreds or thousands of NFE: DDPM requires 1000 NFE, DDIM requires 100 NFE, and Score SDE requires 2000 NFE to achieve their reported FID scores. In contrast, our characteristic generator achieves competitive performance with substantially fewer NFE. For instance, CG achieves FID of 4.59 with only NFE=1, and FID of 2.83 with NFE=4, which is already comparable to Score SDE (FID=2.20) that requires 2000 NFE. This demonstrates that our approach enables significantly faster inference by reducing the number of function evaluations by orders of magnitude. Thus, the offline training cost is an one-time investment that enables efficient inference once and for all.
Appendix N Comparison with Consistency Trajectory Models
In this section, we discuss the implementation details compared with the Consistency Trajectory Models (CTM) [Kim et al., 2024].
N.1. Similarities
  • Our work and the CTM share similar neural network initialization strategy (from a pre-trained teacher model, if any).
  • On MNIST and CIFAR-10, where the teacher and student model are two different networks, our work and CTM shares similar strategies for generating reference solutions from the teacher model: running some ODE solver on the given time interval, with a threshold on the maximum number of steps allowed.
  • Also on MNIST and CIFAR-10, we and CTM both use the Learned Perceptual Image Patch Similarity (LPIPS) metric instead of $L^{2}$-norm to measure distance in the feature space instead of in the original pixel space [Zhang et al., 2018]. We also share similar loss balance techniques to balance between the global loss and the local loss.
N.2. Differences
  • We extend the distillation framework to the latent space on CelebAHQ-256 and CelebA-512, significantly reducing the training and generating cost, while CTM only experiments on the original pixel space. We also manage to shift from the U-Net architecture to the DiT architecture.
  • On CelebAHQ, We extend the distillation framework to self-distillation, and offer a candidate design for the self-guided teacher model, which takes only 2 NFE to generate a reference solution, while CTM still needs dozens of NFEs. Under this setting, the training process of our proposed method would be significantly faster than CTM.
  • The incorporation of GAN contributes a lot to the low FID of CTM, but we find it introduces extra training instability and requires some manual tuning. Our work is more like a clean and minimal template for the distillation-based diffusion models, and the experiments results indicates that our implementation works well enough and could be easily adapted to downstream tasks. We believe that less is more when it comes to high-level framework.
Appendix O Comparisons Between One-Step Euler and Characteristic Generator
  • Error bound for one-step Euler generation. Recall the error bound for the Euler sampling in Theorem III.11:
    $\text{Error of Euler}\leq C\kappa^{2}(T)\Big\{n^{-\frac{2}{d+3}}\log^{2}n+\frac{\log n}{K_{\mathrm{Euler}}^{2}}\Big\},$
    where $K_{\mathrm{Euler}}$ is the number of step of the Euler sampling. One-step Euler generation means $K_{\mathrm{Euler}}=1$, thus
    $\text{Error of one-step Euler}\leq C\kappa^{2}(T)\Big\{n^{-\frac{2}{d+3}}\log^{2}n+\log n\Big\}.$
    The error is dominated by the large $\log n$ term, and does not converge even for sufficiently large number of samples $n$. This large inherent error from using a single, coarse step explains the poor image quality seen in Figure 4.
  • Error bound for one-step characteristic generation. The characteristic generator is fundamentally different. It is a one-step model that is trained to approximate the result of a high-quality, multi-step numerical sampler. In the characteristic learning, the generator are learned from the trajectories generated using an Euler sampler with $K_{\mathrm{CG}}$ steps. Recall the error bound for the characteristic generator in Theorem III.13:
    $\text{Error of CG}\leq C\kappa^{2}(T)\Big\{n^{-\frac{2}{d+3}}\log^{2}n+\frac{\log n}{K_{\mathrm{CG}}^{2}}\Big\}+C\Big\{m^{-\frac{2}{d+4}}\log^{2}m+\frac{\log m}{K_{\mathrm{CG}}}\Big\}.$
    In practical, we can choose sufficiently large $K_{\mathrm{CG}}$. As a result, the final error of the one-step CG is dominated by the sample complexity terms:
    $\text{Error of CG}\leq C\kappa^{2}(T)n^{-\frac{2}{d+3}}\log^{2}n+Cm^{-\frac{2}{d+4}}\log^{2}m,$
    which converge to zero as the dataset sizes ($n$ and $m$) increase.
In essence, the CG distills the knowledge of a precise, multi-step solver into a fast, single-step network. Our theory correctly predicts that its error should be much lower than that of a naive one-step Euler sampler, which aligns perfectly with our empirical findings in Figure 4.